OpenStudy (anonymous):

Look at the pattern below. 324, 108, 36, 12, . . . If this pattern continues, what is the next term? A 6 B 4 C 3 D 1

OpenStudy (destinymasha):

each term is the previous one divided by 3

OpenStudy (destinymasha):

B -4

OpenStudy (anonymous):

Also, wrong subject. I don't understand why you would come here; the Mathematics section of Openstudy is much more populated.

OpenStudy (destinymasha):

sorry if i didnt word it right, each number is divided by 3

OpenStudy (anonymous):

I see that you are in High School based on your profile. If you're in algebra or algebra 2, you should know that (a2)/(a1)=r where r is the common ratio. So 108/324=1/3 Then, use the an=a1r^(n-1) formula to find the next term. You want to find a5 or the 5th item in the sequence, so n=5. You already know that the first thing in the sequence is 324 so plug everything in. a5=324(1/3)^(5-1) IF YOU HAVE NOT LEARNED THIS, IGNORE IT.